Direct Labor Variances

Bellingham Company produces a product that requires 5 standard direct labor hours per unit at a standard hourly rate of $9.00 per hour. If 4,100 units used 19,700 hours at an hourly rate of $9.18 per hour, what is the direct labor (a) rate variance, (b) time variance, and (c) cost variance? Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

a. Direct labor rate variance $
b. Direct labor time variance $
c. Direct labor cost variance $
Answer #1

Labor rate variance = (SR-AR) *AH

= (9-9.18)*19700

= 3546 Unfavorable

Labor time variance = (SH-AH) *SR

= (4100*5 - 19,700)* 9

= -7200 favorable (minus sign)

Labor cost variance = 3546 U + 7200 F

= - 3654 Favourable (minus)